# Quantum resolution limit of long-baseline imaging using distributed entanglement

### v1

@@ VII Generalization to multiple telescopes and more complex imaging problems
Since the QFI-attaining measurement for any single-parameter estimation task—no matter how complex the scene and its configuration—is a von Neumann projective measurement, given the one-photon-per-temporal-conherence-interval assumption, the QFI attaining measurement can always be expressed as a linear interferometric unitary transformation among all the locally-collected spatial modes across all the telescopes in the baseline, followed by shot-noise-limited photon detection on the output modes. We now discuss how the protocol we present in this paper can be generalized to measuring the incoming photons in such an arbitrary linear combination basis of the local modes collected by any number of telescopes. Say we have $n$ telescopes and we are collecting $K$ mutually-orthogonal spatial modes at each site. And our goal is to measure the photons in some arbitrary set of mutually orthogonal linear combinations of these $nK$ signals. If we are carrying out such a measurement by bringing all the $nK$ mode signals to a central location through single mode fibers, then we can use a beam splitter circuit to create an interferometer that mixes all the signals in the desired set of linear combinations. It has been shown in Clements et al. (2016) that we can create such an interferometer that gives us any desired unitary transformation between the $nK$ inputs and the same number of outputs by employing $nK(nK-1)/2$ Mach Zehnder interferometers (MZIs). Since an MZI comprises of two 50-50 beam splitters and two tunable phase shifts, this means a total of $nK(nK-1)$ 50-50 beam splitters and the same number of phase shifts is needed. If we are using pre-shared entanglement to accomplish this losslessly, instead of using (lossy) single mode fibers and beam splitters, then we need three ingredients:
@@
    An entanglement based protocol that mimics a fifty-fifty beam splitter between two modes whose quantum state resides in the span of one photon spread across the two modes, at two distant sites. We will shortly make this statement more precise and show that we can indeed accomplish this.
@@
  3. 3.
    (Optional) An $n$ site generalization of our procedure of using Bell states ancillas to determine the spatio-temporal mode of an incoming photon, to possibly realize the nonlocal version of the Clements et al. circuit Clements et al. (2016) more resource-efficiently. It is straightforward to see that this can be accomplished by employing a GHZ state in place of a Bell state (to realize a three-qubit nonlocal Clifford unitary) and using $CZ$ gates just as we did in Section V.
We now show how to mimic a fifty-fifty beam splitter on the space limited to zero or one photon among two modes. Consider the input modes $A$ and $B$ of a 50-50 beam splitter with outputs $A^{\prime}$ and $B^{\prime}$. The beam splitter will have the action:
@@
$\frac{a+b}{\sqrt{2}}|1_{A^{\prime}}0_{B^{\prime}}\rangle$ $+\frac{a-b}{\sqrt{2}}|0_{A^{\prime}}1_{B^{\prime}}\rangle$ (82)
$\downarrow$
$|0_{A}$ $0_{B}\rangle$ (83)
$\downarrow$
$|0_{A^{\prime}}$ $0_{B^{\prime}}\rangle.$ (84)
Interpreting the above states of two modes as states of two single-rail qubits (i.e., qubits encoded by the absence or the presence of a single photon in a mode), it is easy to see that the above can be accomplished by carrying out the following sequence of operations:
@@
    A CNOT gate from qubit $A$ to $B$.
@@
    The Hadamard gate on qubit $A$.
@@
    A CNOT gate from qubit $A$ to $B$.
With the two qubits at two different sites, the CNOT operations need to be carried out through gate teleportation, which consumes one pre-shared two-qubit Bell state, each, in a way identical to that in Ref.  Gottesman and Chuang (1999). We have thus shown how to mimic a 50-50 beam splitter across two sites. We can therefore use this and apply the framework in Clements et al. (2016) to mimic any nonlocal multimode beam splitter circuit to measure the incoming photons in any arbitrary collective basis.
@@ IX Acknowledgements
The authors thank Michael Grace for helpful discussions, and Prajit Dhara for comments on the manuscript. This work was co-funded by AFOSR grant number FA9550-22-1-0180 and NASA grant number 80NSSC22K1030.

### v2

@@ VII Generalization to multiple telescopes and more complex imaging problems
Since the QFI-attaining measurement for any single-parameter estimation task—no matter how complex the scene and its configuration—is a von Neumann projective measurement, given the one-photon-per-temporal-conherence-interval assumption, the QFI attaining measurement can always be expressed as a linear interferometric unitary transformation among all the locally-collected spatial modes across all the telescopes in the baseline, followed by shot-noise-limited photon detection on the output modes. We now discuss how the protocol we present in this paper can be generalized to measuring the incoming photons in such an arbitrary linear combination basis of the local modes collected by any number of telescopes. Say we have $n$ telescopes and we are collecting $K$ mutually-orthogonal spatial modes at each site. And our goal is to measure the photons in some arbitrary set of mutually orthogonal linear combinations of these $nK$ signals. If we are carrying out such a measurement by bringing all the $nK$ mode signals to a central location through single mode fibers, then we can use a beam splitter circuit to create an interferometer that mixes all the signals in the desired set of linear combinations. It has been shown in Clements et al. (2016) that we can create such an interferometer that gives us any desired unitary transformation between the $nK$ inputs and the same number of outputs by employing $nK(nK-1)/2$ Mach Zehnder interferometers (MZIs). Since an MZI comprises of two 50-50 beam splitters and two tunable phase shifts, this means a total of $nK(nK-1)$ 50-50 beam splitters and the same number of phase shifts is needed. If we are using pre-shared entanglement to accomplish this losslessly, instead of using (lossy) single mode fibers and beam splitters, then we need the following ingredients:
@@
    A general version of interpretting the parities of the $X$ basis measurement results of the photonic qubits and how they translate into relative signs between the various terms in the quantum state. In our two telescope example, where we are only mimicking the classical interferometer that mixes the signals from the two sites associated with the same spatial mode, we only have to consider the parity of the $X$ basis results for a single spatial mode at both sites. In the most general case where we also mix multiple spatial modes together, we will have to consider the $X$ basis results of several spatial modes at various sites simultaneously. Therefore, instead of just a single sign $f(\textbf{h}_{mq})$, we will have to keep track of several such signs in a slightly elaborate (but easily doable) book keeping exercise.
@@
    $n$-site and $K$-spatial mode generalization of our heralding procedure in the decoding section for determining the spatio-temporal mode of an incoming photon with ancilla Bell states, once the state of the incoming photons has been loaded on to quantum memories. It is straightforward to see that using $CZ$ gates in the same way but employing a GHZ state instead of a Bell pair can easily accomplish this. We will shortly describe this more explicitly.
  3. 3.
    An operation on a pair of logical qubits comprising several atomic memories, that mimics the action of a fifty-fifty beam splitter between two optical modes whose quantum state resides in the span of one photon spread across the two modes. These logical qubits may be at the same location or at two distant sites, in which case accomplishing this requires an entangled state shared across the two locations. We will shortly make this statement more precise and show that we can indeed accomplish this.
  4. 4.
@@
  5. 5.
    (Optional) An $n$ site and multiple spatial mode generalization of our procedure of using Bell states ancillas to realize the nonlocal version of the Clements et al. circuit Clements et al. (2016) more resource-efficiently by employing multi-partite entangled resources instead of only using pairwise Bell states and mimicking fifty-fifty beam splitters.
We now clarify the generalization of the heralding procedure with GHZ states. Our heralding procedure with Bell states in the decoding section is based on the fact that if we have a Bell state $\frac{1}{\sqrt{2}}[|00\rangle+|11\rangle]$, then applying a $Z$ gate on one of the two qubits imparts a minus sign on the second term, flipping the Bell state to $\frac{1}{\sqrt{2}}[|00\rangle-|11\rangle]$. Therefore, if we have a pair of memory qubits in the $|10\rangle$ or $|01\rangle$, then applying a $CZ$ gate from one of these memories to one qubit in the Bell pair, and another $CZ$ from the other memory qubit to the second qubit in the Bell pair, flips the sign of the Bell state. But if we have $|00\rangle$ in the two memories, the Bell state will remain unflipped.
This logic clearly also holds when we have a GHZ state. Consider the states
$|GHZ^{\pm}_{n}\rangle\equiv\frac{1}{\sqrt{2}}[|0_{1^{\prime}}0_{2^{\prime}}\ldots 0_{n^{\prime}}\rangle\pm|1_{1^{\prime}}1_{2^{\prime}}\ldots 1_{n^{\prime}}\rangle].$ (82)
Then
$V_{11^{\prime}}|1_{1}0_{2}0_{3}\ldots 0_{n}\rangle|GHZ^{+}_{n}\rangle$ $=$ $V_{22^{\prime}}|1_{1}0_{2}0_{3}\ldots 0_{n}\rangle|GHZ^{+}_{n}\rangle$ (83)
$=$ $\ldots$
$=$ $V_{nn^{\prime}}|0_{1}0_{2}\ldots 0_{n-1}1_{n}\rangle|GHZ^{+}_{n}\rangle$
$=$ $|GHZ^{-}_{n}\rangle,$
where $V_{j,j^{\prime}}$ is a $CZ$ between qubits $j$ and $j^{\prime}$. We can thus generalize our heralding procedure by employing GHZ states.
Also, in our two-telescope example, we determine the temporal and spatial mode simultaneously since we are only measuring in linear combinations involving the same spatial mode collected at both sites. In a general receiver where we are measuring in a basis that mixes two or more spatial modes, we may have a heralding procedure that only determines the temporal mode through the heralding step, while leaving the spatial mode completely undetermined, or only narrows down to certain combinations of the spatial modes. For example, say we have 4 telescopes with $K=4$. If for each $q$, we wish to measure the incoming photon from the scene in some linear combination of the mode $q$ state associated with each site, then we are not mixing the different spatial modes with each other. In that case, a heralding procedure with a 4-component GHZ state for each set of memory qubits (corresponding to a given digit in the binary representation of the temporal modes) can be used to simultaneously determine the temporal and spatial mode just as in our pairwise receiver for two telescopes. However, if we wish to measure the incoming photons in some combination of the different spatial modes, we will not want to employ a heralding procedure that collapses tthe state on to a single spatial mode number $q$. In such a situation, we will design the heralding procedure according to our measurement basis. If there is no simplification where the heralding procedure can help us narrow down to a subspace spanned by a subset of the states in the measurement basis, the most general, brute-force approach would involve a $Kn$-component GHZ state that only determines the temporal mode while keeping the entire spatial mode related structure of the quantum state in tact.
We now show how to mimic a fifty-fifty beam splitter on the Hilbert space spanned by one photon across two modes, $\left\{|1_{A}0_{B}\rangle,|0_{A}1_{B}\rangle\right\}$. Consider the input modes $A$ and $B$ of a 50-50 beam splitter with output modes $A^{\prime}$ and $B^{\prime}$. The beam splitter will have the following action:
@@
$\frac{a+b}{\sqrt{2}}|1_{A^{\prime}}0_{B^{\prime}}\rangle$ $+\frac{a-b}{\sqrt{2}}|0_{A^{\prime}}1_{B^{\prime}}\rangle.$ (85)
In our entanglement-based receiver, we load the photonic state on to quantum memories. With the logarithmic compression of Khabiboulline et al., a (single-rail) photonic qubit for each spatial mode at each site associated with temporal mode $m$ is loaded on to $n_{m}$ (i.e. the number of $1$s in the binary code of $m$ at the same location. If there is no incoming photon in the said spatial mode at that site in temporal mode $m$, then these $n_{m}$ memory qubits remain in $|0\rangle$, and if there is a photon, then all of them are flipped to $|1\rangle$ in the loading process. These $n_{m}$ memory qubits therefore effectively serve as logical qubits associated with the corresponding spatial mode. To obtain the quantum version of the Clements et al. circuit Clements et al. (2016), we need to apply an operation that gives the operation (85) on two such logical qubits, wich may be at the same location or two distant sites. If the states of the input optical modes $A$ and $B$ are transfered on to logical qubits $\overline{E}$ and $\overline{F}$, then the above (beam splitter equivalent) operation can be accomplished by carrying out the following sequence
@@
    A CNOT gate from logical qubit $\overline{E}$ to $\overline{F}$.
@@
    The Hadamard gate on logical qubit $\overline{E}$.
@@
    A CNOT gate from logical qubit $\overline{E}$ to $\overline{F}$.
When the two qubits are at two different sites, the CNOT operations need to be carried out through gate teleportation, which consumes one pre-shared two-qubit Bell state, each, in a way identical to that in Ref.  Gottesman and Chuang (1999). We have thus shown how to mimic a 50-50 beam splitter across two sites. We can therefore use this and apply the framework in Clements et al. (2016) to mimic any nonlocal multimode beam splitter circuit to measure the incoming collected optical field across $n$ remotely-located modes in any arbitrary collective basis. This is true as long as the quantum state of those modes lives in the unary subspace, i.e., one photon spread across the $n$ modes. This condition is satisfied as part of the premise of the problem setup as described above.
@@ IX Acknowledgements
The authors thank Michael Grace, Gabe Richardson, Brittany McClinton, Jayadev Rajagopal, Ryan Lau, and Stephen Ridgway for helpful discussions, and Prajit Dhara for comments on the manuscript. This work was co-funded by AFOSR grant number FA9550-22-1-0180 and NASA grant number 80NSSC22K1030.